- E-beam is finished. The staff is using it to catch up on foundry work and it will be available Monday for users. We still need to dissemble and clean the emitter and we plan to purchase a replacement so we have one in stock to minimize down time.
- Thermal evaporator is back online. Bill Tang’s group will be using it to deposit Ti.
- We are planning to address the chill water system with the STS starting Monday. This will impact the Thermal Evaporator, STS and Plasma Therm. To revamp the existing chill water it will take five working days.
- The replacement lamp for the Flood Exposure in the BiON came in yesterday and will be replaced this afternoon. The system should be up and running later today.